# Multiscale Hodge Scattering Networks for Data Analysis

### v5

@@ 4 Theoretical Analysis of Multiscale Hodge Scattering Transforms
In this section we establish approximation properties of the multiscale basis dictionaries, then use there results to detail the continuity, invariance and equivalence properties of the MHSNs which make them desirable for signal and domain classification problems. For notational convenience, we let $S=S^{1}$ and only consider the first order transform in our formal proofs. However, since $S^{m}{\boldsymbol{f}}$ can be thought of as applying $S^{1}$ to the transform $S_{0}^{m-1}{\boldsymbol{f}}$, all of the proofs can be trivially generalized to the $S^{m}$ case.
First, for singleton $\kappa$-elements $\sigma$ and $\tau$ of $C_{\kappa}$, and signal ${\boldsymbol{f}}$, we define a distance function and then the associated Hölder semi-norm as the number of elements in the smallest partition in the tree that constrains both elements, formally we have:
$d(\sigma,\tau)\,:=\,\min\left\{n^{j}_{k}\,\Big{|}\,\sigma,\tau\in C^{j}_{k}\right\},\quad C_{H}({\boldsymbol{f}})\,:=\,\sup_{\sigma\neq\tau}\frac{\left|[{\boldsymbol{f}}]_{\sigma}-[{\boldsymbol{f}}]_{\tau}\right|}{d(\sigma,\tau)^{\alpha}}$
where $\alpha$ is a constant in $(0,1]$. With these definitions, the dictionary coefficient decay and approximation results of [32, 6] for the GHWT and HGLET can be applied to the $\kappa$-GHWT and $\kappa$-HGLET bases as detailed further in [7].
Theorem 1 (Fast Decay of Multiscale Coefficients).
For a simplicial complex $C$ equipped with a hierarchical bipartition tree, suppose that a signal ${\boldsymbol{f}}$ is Hölder continuous with exponent $\alpha$ and constant $C_{H}({\boldsymbol{f}})$. Then the coefficients with $l\geq 1$ for the $\kappa$-HGLET ($c^{j}_{k,l}$) and $\kappa$-GHWT ($d^{j}_{k,l}$) satisfy:
$\left|c^{j}_{k,l}\right|\leq C_{H}({\boldsymbol{f}})\left(n^{j}_{k}\right)^{\alpha+\frac{1}{2}},\quad\left|d^{j}_{k,l}\right|\leq C_{H}({\boldsymbol{f}})\left(n^{j}_{k}\right)^{\alpha+\frac{1}{2}}$
@@
See Theorem 3.1 of [6]. ∎
Theorem 2 (Approximation Power of Multiscale Bases).
For a fixed ONB $\{{\boldsymbol{\phi}}_{l}\}^{n-1}_{l=0}$ and a parameter $20<\rho<$ 2, then
$\|{\boldsymbol{f}}-P_{m}{\boldsymbol{f}}\|_{2}\leq\frac{|{\boldsymbol{f}}|_{\rho}}{m^{\beta}}.$
where $P_{m}$ in the best nonlinear $m$-term approximation in the basis $\{{\boldsymbol{\phi}}_{l}\}^{n-1}_{l=0}$, $\beta=\frac{1}{\rho}-\frac{1}{2}$ and $|{\boldsymbol{f}}|_{\rho}$ is defined as $|{\boldsymbol{f}}|_{\rho}\,:=\,\left(\sum_{l=0}^{n-1}|\langle{\boldsymbol{f}},{\boldsymbol{\phi}}_{l}\rangle|^{\rho}\right)^{\frac{1}{\rho}}$
@@
See Theorem 3.2 of [6] and Theorem 6.3 of [32]. ∎
Next we establish bounds on the coefficients generate by the multiscale basis. Let ${\mathscr{W}}$ indicate the matrix formed by stacking a multiscale basis into a $(J+1)n\times n$ matrix where the $j$th $n\times n$ block is the $j$th-level ONB. Next we define a weighted inner product and the associated norm for signals in $X_{\kappa}$ as:
$\langle{\boldsymbol{f}},\boldsymbol{g}\rangle_{w}\,:=\,\sum_{\sigma\in C_{\kappa}}[{\boldsymbol{f}}]_{\sigma}[{\boldsymbol{g}}]_{\sigma}[{\boldsymbol{w}}]_{\sigma}\quad\text{and}\quad\|{\boldsymbol{f}}\|_{w}\,:=\,\sqrt{\langle{\boldsymbol{f}},{\boldsymbol{f}}\rangle_{w}}.$
In practice, we can often use ${\boldsymbol{w}}\equiv{\boldsymbol{1}}$, but for some applications more exotic norms, such as those that consider the volume of each face, may be useful.
Remark 1 (Tight Frames).
Let ${\boldsymbol{f}}\in X_{\kappa}$, and $\|\cdot\|_{w}$ indicate the $\ell^{2}$-norm with respect to the metric $w$. Then we have:
$\|{\boldsymbol{f}}\|^{2}_{w}\leq\|{\mathscr{W}}{\boldsymbol{f}}\|_{w}^{2}=(J+1)\|{\boldsymbol{f}}\|^{2}_{w}$
This remark is clearly true since ${\mathscr{W}}$ is simply a collection of $J+1$ orthogonal matrices, and orthogonal transforms preserve the $\ell^{2}$-norm. Although trivial, this fact will be vital for later proofs. Next we show that the MHST is a non-expansive operation. This allows for powerful nonlinear analysis properties [33] which we detail later.
Theorem 3 (Non-expansive Operation).
Let $S$ be the MHST with global pooling formed by the multiscale basis dictionary ${\mathscr{W}}$ as defined above in (3)–(5), acting on the metric space ${\mathscr{W}}$. Then we have:
$\|S{\boldsymbol{f}}_{1}-S{\boldsymbol{f}}_{2}\|_{w}\leq\|{\boldsymbol{f}}_{1}-{\boldsymbol{f}}_{2}\|_{w}$
[…]
$\|S_{j}{\boldsymbol{f}}_{1}-S_{j}{\boldsymbol{f}}_{2}\|_{w}\leq\|{\boldsymbol{f}}_{1}-{\boldsymbol{f}}_{2}\|_{w}$
Proof.
To show that the MHST is non-expansive, we will show that each layer is non-expansive. Since the entire transform is defined by a cascade of these layers (with modulus operations taken between the layers), it will also be non-expansive. Then:
$S^{1}\left(q,j\right)({\boldsymbol{f}}_{1}-{\boldsymbol{f}}_{2})$ $=\frac{1}{n}\sum_{i=1}^{n}\left|[\Phi^{j}({\boldsymbol{f}}_{1}-{\boldsymbol{f}}_{2})]_{i}\right|^{q}$ (9)
$\leq\left|\frac{1}{n}\sum_{i=1}^{n}\left|[\Phi^{j}{\boldsymbol{f}}_{1}]_{i}\right|^{q}-\frac{1}{n}\sum_{i=1}^{n}\left|[\Phi^{j}{\boldsymbol{f}}_{2}]_{i}\right|^{q}\right|=\left|S^{1}\left(q,j\right){\boldsymbol{f}}_{1}-S^{1}\left(q,j\right){\boldsymbol{f}}_{2}\right|$
Since this inequality holds for each $q$ and $j$, it is clear that taking the $w$-norm over the whole collection of features $S^{1}{\boldsymbol{f}}$ will also hold. Then, since each order of the transform features can be formed by applying $S^{1}$ to the previous order we have: $\|S{\boldsymbol{f}}_{1}-S{\boldsymbol{f}}_{2}\|_{w}\leq\|{\boldsymbol{f}}_{1}-{\boldsymbol{f}}_{2}\|_{w}$. The proof for the local summation follows exactly as above. ∎
Next we show that our networks are invariant under group operations which preserve the inner product $\langle\cdot,\cdot,\rangle_{w}$, such as permutations of the $\kappa$-simplices. That is, relabeling indices of the elements of $C_{\kappa}$ does not affect the output of the transform. For example, given a signal ${\boldsymbol{f}}$ defined on the triangular faces of a simplex, the indexing of the triangles does not effect the globally pooled transform and permuting this indexing results in an equivalent permutation of the non-pooled signal. This theorem is analogous to Theorem 3 in [25] and Proposition 4.1 in [34], but generalizes to any $\kappa$, rather than strictly to the nodes ($\kappa$=0 case).
Theorem 4 (Invariance and Equivariance).
Suppose $\mathcal{G}$ is a group of transforms on the elements of $C_{\kappa}$(e.g., permutations of the $\kappa$-simplices). Furthermore, given any ${\boldsymbol{\xi}}\in\mathcal{G}$, let $V_{\boldsymbol{\xi}}:C_{\kappa}\rightarrow C_{\kappa}$ denote the operator $V_{\boldsymbol{\xi}}{\boldsymbol{f}}={\boldsymbol{f}}\circ{\boldsymbol{\xi}}$ induced by ${\boldsymbol{\xi}}$ and $S^{[{\boldsymbol{\xi}}]}$ the analogous transform on $C^{[{\boldsymbol{\xi}}]}_{\kappa}$, which is the permuted version of $C_{\kappa}$. Then both of the following hold:
$S^{[{\boldsymbol{\xi}}]}V_{\boldsymbol{\xi}}{\boldsymbol{f}}=S{\boldsymbol{f}}\quad\text{ and }\quad S_{j}^{[{\boldsymbol{\xi}}]}V_{\boldsymbol{\xi}}{\boldsymbol{f}}=S_{j}{\boldsymbol{f}}$ (10)
for all ${\boldsymbol{\xi}}\in\mathcal{G}$, ${\boldsymbol{f}}\in{\mathbb{R}}^{n}$ and $0\leq j\leq J$.
Proof.
As with the previous proof we will show that this holds for an arbitrary layer, and since the entire transform is formed from a cascade of these layers, it will also have this property. First, denote $\mathcal{L}_{\boldsymbol{\xi}}=V_{\boldsymbol{\xi}}\circ\mathcal{L}\circ V_{\boldsymbol{\xi}}^{-1}$ where $\mathcal{L}$ is an appropriate version of the $\kappa$-Laplacian for $C_{\kappa}$. It is immediately clear that if ${\boldsymbol{\phi}}$ is an eigenvector of $\mathcal{L}$ with $\mathcal{L}{\boldsymbol{\phi}}=\lambda{\boldsymbol{\phi}}$ then $V_{\boldsymbol{\xi}}{\boldsymbol{\phi}}$ is an eigenvector of $\mathcal{L}_{\boldsymbol{\xi}}$, i.e., $\mathcal{L}_{\boldsymbol{\xi}}V_{\boldsymbol{\xi}}{\boldsymbol{\phi}}=\lambda V_{\boldsymbol{\xi}}{\boldsymbol{\phi}}$ since
$\mathcal{L}_{\boldsymbol{\xi}}V_{\boldsymbol{\xi}}=V_{\boldsymbol{\xi}}\circ\mathcal{L}\circ V_{\boldsymbol{\xi}}^{-1}\circ V_{\boldsymbol{\xi}}=V_{\boldsymbol{\xi}}\circ\mathcal{L}$ (11)
@@
Then Since $\mathcal{G}$ preserves the inner product on ${\mathscr{W}}$, it is clear that $V_{\boldsymbol{\xi}}\Phi\,:=\,\Phi_{\boldsymbol{\xi}}$ is an ONB for ${\mathscr{W}}_{\xi}$ whose atoms are the bases $\Phi_{\xi}$ formed from eigenvectors of $\mathcal{L}_{\boldsymbol{\xi}}$, where $\Phi_{\xi}=V_{\xi}\Phi V_{\xi^{-1}}$
Then for arbitrary input $f$ we have:
$S^{[{\boldsymbol{\xi}}]}V_{\boldsymbol{\xi}}{\boldsymbol{f}}=\frac{1}{n}\sum_{i=1}^{n}\left|[\Phi_{\boldsymbol{\xi}}V_{\boldsymbol{\xi}}{\boldsymbol{f}}]_{i}\right|=\frac{1}{n}\sum_{i=1}^{n}\left|[V_{\boldsymbol{\xi}}\Phi{\boldsymbol{f}}]_{i}\right|=\frac{1}{n}\sum_{i=1}^{n}\left|[\Phi{\boldsymbol{f}}]_{i}\right|=S{\boldsymbol{f}}$ (12)
The proof for the local summation follows exactly as above. ∎
@@ 5 Signal Classification
We first demonstrate the effectiveness of our MHSNs with the article category classification problem using the Science News database [35, 36]. We do not claim state-of-the-art results for this problem, but rather use it to illustrate the advantages of analyzing signals via higher-dimensional simplices. We also show how locally-pooled networks can shatter problems in which traditional globally-pooled scattering networks fail to differentiate between classes. The Science News dataset contains 1042 scientific news articles classified into eight fields: Anthropology, Astronomy; Behavioral Sciences; Earth Sciences; Life Sciences; Math/CS; Medicine; Physics. Each article is tagged with keywords from a pool of 1133 words selected by the database curators. We determine a simplicial complex from these keywords by computing their word2vec [37] embeddings based on Google’s publicly available pre-trained model [38]. We generate a symmetric $k$-nearest neighbor graph of the embedded words and then generate $\kappa$-simplices of the graph. Therefore, a $\kappa$-simplex in this keyword graph corresponds to $\kappa$-face, which represents a combination of $\kappa+1$ words.
Next, we define representations of each article as a signal in each $X_{\kappa}$ as follows. First, for $\kappa=0$ (i.e., a node-valued signal), we define the signal ${\boldsymbol{f}}_{0}$ to be one on the nodes representing their keywords and zero elsewhere. For $\kappa\geq 1$ we define the signal ${\boldsymbol{f}}_{\kappa}$ to be the simplex-wise average of the ${\boldsymbol{f}}_{0}$ signal. That is,
$[{\boldsymbol{f}}_{0}]_{i}=\begin{cases}1&\text{if keyword $i$ occurs}\\ 0&\text{Otherwise}\end{cases};\quad[{\boldsymbol{f}}_{\kappa}]_{i}=\frac{1}{\kappa+1}\sum_{\begin{subarray}{c}l\in V(\sigma_{i})\\ \sigma_{i}\in C_{\kappa}\end{subarray}}[{\boldsymbol{f}}_{0}]_{l},$ (13)
where $V(\sigma_{i})$ represents the set of nodes forming the $i$th simplex $\sigma_{i}\in C_{\kappa}$. Note that these signals are highly localized since the keywords are connected through a symmetrized $k$NN graph, and the higher-order signals are built from the adjacency of the resulting complex. To showcase the robustness of our approach, we report results using both $k=5$ and $k=10$ nearest neighbor graphs.
@@
Table 1: Article category classification accuracy for $5$-NN graph of the Science News dataset for different simplex degrees. Dict. implies that the SVM is trained solely on the dictionary coefficients while GP, LP, NP imply scattering networks with global, local, and no pooling, respectively. The best performer for each $\kappa$ is indicated in bold while the underlined bold numbers are the best among all $\kappa$’s.
@@
Table 2: Article category classification accuracy for $10$-NN graph of the Science News dataset for different simplex degrees. Dict. implies that the SVM is trained solely on the dictionary coefficients while GP, LP, NP imply scattering networks with global, local, and no pooling, respectively. The best performer for each $\kappa$ is indicated in bold while the underlined bold numbers are the best among all $\kappa$’s.

### v6

@@ 4 Theoretical Analysis of Multiscale Hodge Scattering Transforms
In this section, we establish the continuity, invariance, and equivalence properties of the MHSNs that make them well suited for signal and domain classification tasks. For notational convenience, we let $S\equiv S^{1}$ and restrict our formal proofs to the first-order transform. However, since $S^{m}\bm{f}$ can be viewed as applying $S^{1}$ to the transform $S_{0}^{m-1}\bm{f}$, followed by the pooling operation described in Eq. (4), all of the proofs extend directly to the general $S^{m}$ case.
First, we establish bounds on the coefficients generated by the multiscale basis. Let $\mathcal{W}$ denote the matrix obtained by stacking a multiscale basis into a $(J+1)n\times n$ matrix, where the $j$th $n\times n$ block corresponds to the ONB at scale $j$. We define a weighted inner product and the associated norm for signals in $X_{\kappa}$ by
$\langle\bm{f},\bm{g}\rangle_{w}\,:=\,\sum_{\sigma\in C_{\kappa}}[\bm{f}]_{\sigma}[\bm{g}]_{\sigma}[\bm{w}]_{\sigma}\quad\text{and}\quad\|\bm{f}\|_{w}\,:=\,\sqrt{\langle\bm{f},\bm{f}\rangle_{w}}.$
In practice, we often take $\bm{w}\equiv{\bm{1}}$, but in some applications more specialized weightings—such as those incorporating the volume of each face—may be advantageous.
Remark 1 (Tight Frames).
Let $\bm{f}\in X_{\kappa}$, and $\|\cdot\|_{w}$ denote the $\ell^{2}$-norm with respect to the weight vector $\bm{w}$. Then,
$\|\bm{f}\|_{w}^{2}\leq\|\mathcal{W}\bm{f}\|_{w}^{2}=(J+1)\|\bm{f}\|^{2}_{w}\,.$
This follows immediately from the fact that $\mathcal{W}$ is the vertical concatenation of $J+1$ orthogonal matrices, and orthogonal transforms preserve the $\ell^{2}$-norm. Although elementary, this property will play a key role in subsequent proofs. We next show that the MHST is a non-expansive operator, enabling the use of powerful tools from nonlinear contraction analysis [34], as illustrated in our numerical experiments.
Theorem 1 (Non-expansive Operation).
Let $S$ denote the MHST with global pooling, formed from the multiscale basis dictionary $\mathcal{W}$ as defined in (1)–(3), with $q=1$, acting on the metric space $(X_{\kappa},\|\cdot\|_{w})$. Then, for all $\bm{f}_{1},\bm{f}_{2}\in X_{\kappa}$,
$\|S\bm{f}_{1}-S\bm{f}_{2}\|_{w}\leq\|\bm{f}_{1}-\bm{f}_{2}\|_{w}\,.$
Moreover, let $S_{j}$ denote the transforms with local pooling as described in Section 3.1. Then, for all $0\leq j\leq J$ and all $\bm{f}_{1},\bm{f}_{2}\in X_{\kappa}$,
$\|S_{j}\bm{f}_{1}-S_{j}\bm{f}_{2}\|_{w}\leq\|\bm{f}_{1}-\bm{f}_{2}\|_{w}$
@@
We first show that each layer of the MHST is non-expansive. Since the full transform is a cascade of such layers, with pointwise modulus operations between them, the entire transform will also be non-expansive.
Fix $j$ and consider the first-order coefficient $S^{1}(1,j)\bm{f}$. For $\bm{f}_{1},\bm{f}_{2}\in X_{\kappa}$,
$\left|S^{1}\left(1,j\right)\bm{f}_{1}-S^{1}\left(1,j\right)\bm{f}_{2}\right|$ $=\left|\frac{1}{n}\sum_{i=1}^{n}\left[\left|\Phi^{j}\bm{f}_{1}\right|\right]_{i}-\frac{1}{n}\sum_{i=1}^{n}\left[\left|\Phi^{j}\bm{f}_{2}\right|\right]_{i}\right|$
$\leq\frac{1}{n}\sum_{i=1}^{n}\left|\left[\left|\Phi^{j}\bm{f}_{1}\right|\right]_{i}-\left[\left|\Phi^{j}\bm{f}_{2}\right|\right]_{i}\right|$
$\leq\frac{1}{n}\sum_{i=1}^{n}\left|\left[\Phi^{j}\bm{f}_{1}\right]_{i}-\left[\Phi^{j}\bm{f}_{2}\right]_{i}\right|$
$=\frac{1}{n}\sum_{i=1}^{n}\left|\left[\Phi^{j}(\bm{f}_{1}-\bm{f}_{2})\right]_{i}\right|$
$=\frac{1}{n}\sum_{i=1}^{n}\left|\left\langle{\bm{f}_{1}-\bm{f}_{2}},{\Phi^{j}(i,:)}\right\rangle\right|$
$\leq\frac{1}{n}\sum_{i=1}^{n}\left\|\bm{f}_{1}-\bm{f}_{2}\right\|_{w}\left\|\Phi^{j}(i,:)\right\|_{w}\quad\text{(Cauchy--Schwarz)}$
$=\frac{1}{n}\sum_{i=1}^{n}\left\|\bm{f}_{1}-\bm{f}_{2}\right\|_{w}$
$=\|\bm{f}_{1}-\bm{f}_{2}\|_{w}\,$,
where we have used the fact that each $\Phi^{j}$ is an ONB w.r.t. $\|\cdot\|_{w}$, so $\|\Phi^{j}(i,:)\|_{w}=1$ for all $i$. Since this inequality holds for each $j$, taking the $w$-norm over the entire collection of first-order features $S^{1}\bm{f}$ preserves the bound. Moreover, higher‑order features are obtained by applying $S^{1}$ to the previous order’s output (followed by pooling), so the same argument applies inductively. Thus
$\|S\bm{f}_{1}-S\bm{f}_{2}\|_{w}\leq\|\bm{f}_{1}-\bm{f}_{2}\|_{w}\,.$
The proof for the local-pooling transforms $S_{j}$ follows identically, with the global average replaced by the average over each local region. ∎
Next, we show that our networks are invariant under group operations that preserve the weighted inner product $\langle\cdot,\cdot,\rangle_{w}$, such as permutations of the $\kappa$-simplices. In other words, relabeling the indices of the elements of $C_{\kappa}$ does not affect the output of the transform. For example, given a signal $\bm{f}$ defined on the triangular faces of a simplex, the indexing of the triangles does not affect the globally-pooled transform, and permuting this indexing produces the corresponding permutation of the non-pooled signal. This result is analogous to Theorem 3 in [28] and Proposition 4.1 in [8], but it applies to any $\kappa$, rather than only to the node case ($\kappa=0$).
Theorem 2 (Invariance and Equivariance).
Let $\mathcal{G}$ be a group of transformations on the elements of $C_{\kappa}$ (e.g., permutations of the $\kappa$-simplices). For any $\bm{{\xi}}\in\mathcal{G}$, let $V_{\bm{{\xi}}}:X_{\kappa}\to X_{\kappa}$ denote the operator $V_{\bm{{\xi}}}\bm{f}=\bm{f}\circ\bm{{\xi}}$, i.e., $(V_{\bm{{\xi}}}\bm{f})(\sigma)=\bm{f}\left(\bm{{\xi}}(\sigma)\right)$, induced by $\bm{{\xi}}$. Let $S^{[\bm{{\xi}}]}$ denote the analogous transform on $C^{[\bm{{\xi}}]}_{\kappa}$, the permuted version of $C_{\kappa}$. Then, for $\bm{{\xi}}\in\mathcal{G}$, $\bm{f}\in X_{\kappa}$, and $0\leq j\leq J$,
$S^{[\bm{{\xi}}]}V_{\bm{{\xi}}}\bm{f}=S\bm{f}\,,\qquad S_{j}^{[\bm{{\xi}}]}V_{\bm{{\xi}}}\bm{f}=S_{j}\bm{f}\,.$
@@
As in the previous proof, it suffices to show that the property holds for an arbitrary layer; since the full transform is a cascade of such layers, the result then follows for the entire transform.
Let $\mathcal{L}_{\bm{{\xi}}}=V_{\bm{{\xi}}}\circ\mathcal{L}\circ V_{\bm{{\xi}}}^{-1}$, where $\mathcal{L}$ is the $\kappa$-Laplacian on $C_{\kappa}$. If $\bm{{\phi}}$ is an eigenvector of $\mathcal{L}$ with $\mathcal{L}\bm{{\phi}}=\lambda\bm{{\phi}}$, then
$\mathcal{L}_{\bm{{\xi}}}V_{\bm{{\xi}}}\bm{{\phi}}=V_{\bm{{\xi}}}\circ\mathcal{L}\circ V_{\bm{{\xi}}}^{-1}\circ V_{\bm{{\xi}}}\bm{{\phi}}=V_{\bm{{\xi}}}\circ\mathcal{L}\bm{{\phi}}=\lambda\,V_{\bm{{\xi}}}\bm{{\phi}}.$
Thus $V_{\bm{{\xi}}}\bm{{\phi}}$ is an eigenvector of $\mathcal{L}_{\bm{{\xi}}}$ with the same eigenvalue $\lambda$.
@@
Since $\mathcal{G}$ preserves the weighted inner product on $\mathcal{W}$, it follows that $\Phi_{\bm{{\xi}}}\,:=\,V_{\bm{{\xi}}}\Phi$ is an ONB for $\mathcal{W}_{\bm{{\xi}}}$, where the atoms of $\Phi_{\bm{{\xi}}}$ are eigenvectors of $\mathcal{L}_{\bm{{\xi}}}$. In particular, $\Phi_{\bm{{\xi}}}=V_{\bm{{\xi}}}\Phi V_{\bm{{\xi}}^{-1}}$.
Now, for arbitrary input $f\in X_{\kappa}$,
$S^{[\bm{{\xi}}]}V_{\bm{{\xi}}}\bm{f}=\frac{1}{n}\sum_{i=1}^{n}\left|[\Phi_{\bm{{\xi}}}V_{\bm{{\xi}}}\bm{f}]_{i}\right|=\frac{1}{n}\sum_{i=1}^{n}\left|[V_{\bm{{\xi}}}\Phi\bm{f}]_{i}\right|=\frac{1}{n}\sum_{i=1}^{n}\left|[\Phi\bm{f}]_{i}\right|=S\bm{f}$
The proof for the local‑pooling transforms $S_{j}$ is identical, with the global average replaced by the average over each local region. ∎
@@ 5 Signal Classification
We first demonstrate the effectiveness of our MHSNs with the article category classification problem using the Science News database [35, 36]. We do not claim SotA results for this problem, but rather use it to illustrate the advantages of analyzing signals via higher-dimensional simplices. We also show how locally-pooled networks can shatter problems in which traditional globally-pooled scattering networks fail to differentiate between classes. The Science News dataset contains 1042 scientific news articles classified into eight fields: Anthropology, Astronomy; Behavioral Sciences; Earth Sciences; Life Sciences; Math/CS; Medicine; Physics. Each article is tagged with keywords from a pool of 1133 words selected by the database curators. We determine a simplicial complex from these keywords by computing their word2vec [37] embeddings based on Google’s publicly available pre-trained model [38]. We generate a symmetric $k$-nearest neighbor graph of the embedded words and then generate $\kappa$-simplices of the graph. Therefore, a $\kappa$-simplex in this keyword graph corresponds to $\kappa$-face, which represents a combination of $\kappa+1$ words.
Next, we define representations of each article as a signal in each $X_{\kappa}$ as follows. First, for $\kappa=0$ (i.e., a node-valued signal), we define the signal $\bm{f}_{0}$ to be one on the nodes representing their keywords and zero elsewhere. For $\kappa\geq 1$ we define the signal $\bm{f}_{\kappa}$ to be the simplex-wise average of the $\bm{f}_{0}$ signal. That is,
$\bm{f}_{0}]_{i}=\begin{cases}1&\text{if keyword $i$ occurs}\\ 0&\text{otherwise}\end{cases};\quad[\bm{f}_{\kappa}]_{i}=\frac{1}{\kappa+1}\sum_{\begin{subarray}{c}l\in V(\sigma_{i})\\ \sigma_{i}\in C_{\kappa}\end{subarray}}[\bm{f}_{0}]_{l},$
where $V(\sigma_{i})$ represents the set of nodes forming the $i$th simplex $\sigma_{i}\in C_{\kappa}$. Note that these signals are highly localized since the keywords are connected through a symmetrized kkNN graph, and the higher-order signals are built from the adjacency of the resulting complex. To showcase the robustness of our approach, we report results using both $k=5$ and $k=10$ nearest neighbor graphs.
@@
Table 1: Article category classification accuracy for 55-NN graph of the Science News dataset for different simplex degrees. Dict. implies that the SVM is trained solely on the dictionary coefficients while GP, LP, NP imply scattering networks with global, local, and no pooling, respectively. The best performer for each $\kappa$ is indicated in bold while the underlined bold numbers are the best among all $\kappa$’s.
@@
Table 2: Article category classification accuracy for 1010-NN graph of the Science News dataset for different simplex degrees. Dict. implies that the SVM is trained solely on the dictionary coefficients while GP, LP, NP imply scattering networks with global, local, and no pooling, respectively. The best performer for each $\kappa$ is indicated in bold while the underlined bold numbers are the best among all $\kappa$’s.